Circular No. 8635

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2005 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag ≈ 11.5) located at α = 5^{h}10^{m}.6, δ = −69^{o}13′ (equinox 2000.0), which is near the compact cluster NGC 1856, on two unfiltered Technical Pan films taken on Nov. 26.164 UT with a 0.2-m Schmidt camera; the new object is also present at mag ∼ 12.8 on a film taken on Nov. 22.065. Nothing brighter than mag 13 appears at this position on earlier photographs, the most recent taken on Nov. 20. Liller forwards the following precise position measured by B. Allen (Blenheim, New Zealand) from a CCD image taken on Nov. 27.41: α = 5^{h}10^{m}33^{s}.79, δ = −69^{o}12′35′′.2 (V = 12.6; reference star GSC 9161.746) — showing clearly that the new object is not coincident with any star visible on Palomar Sky Survey or Digitized Sky Survey images (limiting mag ∼ 20). Visual magnitude estimate by A. Pearce, Nedlands, W. Australia: Nov. 27.576, 12.8.

F. M. Walter, Stony Brook University; A. Pasten, Cerro Tololo Interamerican Observatory (CTIO); and H. E. Bond, Space Telescope Science Institute, report that a low-resolution spectrogram (range 610–950 nm; resolution 0.86 nm) of Liller’s new object, obtained on Nov. 28.28 UT with the SMARTS/CTIO 1.5-m telescope (+ RC spectrograph), shows strong Hα emission (equivalent width −4.1 nm). O I 777.4- and 814.6-nm lines show P-Cyg profiles, and the Ca II infrared triplet is seen in emission. The new object thus appears to be a classical nova near maximum.

SUPERNOVAE 2005kv–2005kx The discoveries of three apparent supernovae have been reported from

unfiltered CCD frames: 2005kv and 2005kw by the “Nearby Supernova Factory” collaboration from NEAT images, and 2005kx by T. Puckett and J. Tigner from a frame taken with the 0.35-m automated supernova patrol telescope (cf. IAUC 8630) in Ellijay, GA. SNe 2005kv and 2005kw are type-Ia supernovae, ≈ 1 month past maximum and near maximum, respectively (details on CBET 306). Additional approximate magnitudes for SN 2005kx in NGC 3197: Nov. 1, [20.0; 27.00 UT, 17.2 (0.60-m reflector).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005kv Nov. 19.5 8^{h}34^{m}11.^{s}15 + 6^{o}16′34′′.9 19.2  —
2005kw Nov. 19.5 9 36 06.05 +24 24 13.8 20.0  —
2005kx Nov. 26.39 10 14 26.09 +77 49 19.4 17.3 5′′.1 W, 6′′.2 N
2005 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green